OpenStudy (anonymous):

What would be the ordered pair is a solution of y= -5x + 10

hartnn (hartnn):

do u have choices/options to choose from ?

OpenStudy (anonymous):

A.) (-15,5) B.) (2,0) C.) (-1,-5) D.) (3,25)

OpenStudy (texaschic101):

there are several solutions to this...unless, of course, it is multiply choice. NVM.....sub in your answer choices and see which one is true. y = -5x + 10 (2,0) x = 2 and y = 0 sub 0 = -5(2) + 10 0 = -10 + 10 0 = 0 (correct)

OpenStudy (anonymous):

Thank you All very much :)

OpenStudy (texaschic101):

:) happy to help